Exercise 19-3 Income reporting under absorption costing and variable costing LO P2

Sims Company, a manufacturer of tablet computers, began operations on January 1, 2017. Its cost and sales information for this year follows.

Manufacturing costs
Direct materials $ 35 per unit
Direct labor $ 55 per unit
Overhead costs for the year
Variable overhead $ 4,400,000
Fixed overhead $ 6,600,000
Selling and administrative costs for the year
Variable $ 775,000
Fixed $ 4,000,000
Production and sales for the year
Units produced 110,000 units
Units sold 80,000 units
Sales price per unit $ 360 per unit


1.
Prepare an income statement for the year using variable costing.
2. Prepare an income statement for the year using absorption costing.
3. Under what circumstance(s) is reported income identical under both absorption costing and variable costing?
Answer #1
1
SIMS COMPANY
Variable Costing Income Statement
Sales 28800000 =80000*360
Less: Variable costs
Direct materials 2800000 =80000*35
Direct labor 4400000 =80000*55
Variable overhead costs 4400000
Variable selling and administrative expenses 775000
Total variable costs 12375000
Contribution margin 16425000
Less: Fixed expenses
Fixed selling and administrative costs 4000000
Fixed overhead costs 6600000
Total fixed expenses 10600000
Net income (loss) 5825000
2
SIMS COMPANY
Absorption Costing Income Statement
Sales 28800000
Less: Cost of goods sold
Direct materials 2800000
Direct labor 4400000
Variable overhead costs 4400000
Fixed overhead costs 4800000 =6600000/110000*80000
Cost of goods sold 16400000
Gross margin 12400000
Selling general and administrative expenses
Fixed selling and administrative costs 4000000
Variable selling and administrative expenses 775000
Total selling, general and administrative expenses 4775000
Net income (loss) 7625000
3
Production equals sales and there is no beginning finished goods inventory